OpenStudy (isaiah.feynman):

Mathematics puzzle. Express numbers in terms of 2's with a maximum of four 2's.

OpenStudy (isaiah.feynman):

Starting from one.

OpenStudy (anonymous):

1 = 2 - 2 + 2 / 2.

OpenStudy (anonymous):

2 = 2 / 2 + 2 / 2

OpenStudy (anonymous):

3 = (2 + 2 + 2)/2

OpenStudy (anonymous):

4 = 2 + 2 + 2 - 2.

OpenStudy (isaiah.feynman):

And 5? :P

OpenStudy (anonymous):

5 = 2 + 2 + 2 / 2. :)

OpenStudy (isaiah.feynman):

6/2 is not 5 lol

OpenStudy (anonymous):

Nope. It's 2 + 2 + 2 / 2 = 2 + 2 + 1 = 5. :))

OpenStudy (isaiah.feynman):

Oh I see. Nice.

OpenStudy (isaiah.feynman):

gets difficult as the numbers get bigger.

OpenStudy (anonymous):

Yeah. :))

OpenStudy (anonymous):

Spin-off Question: What is the largest integer that can be expressed using 4 2's and any mathematical operation?

OpenStudy (isaiah.feynman):

There is a general formula to express any integer in terms of 2's with a maximum of 4 2's. I dunno it

OpenStudy (isaiah.feynman):

Lets jump to 10!

OpenStudy (anonymous):

10 = 2 x 2 x 2 + 2

OpenStudy (isaiah.feynman):

@iambatman You came and left. Get back here! :P

OpenStudy (isaiah.feynman):

11?

OpenStudy (anonymous):

This one is hard. But, \[11=\frac{(2 + 2)! -2}{2}.\]

OpenStudy (isaiah.feynman):

Ingenious!

OpenStudy (anonymous):

Thanks. :))

OpenStudy (isaiah.feynman):

\[12= 2(2^{2}+2)\]

OpenStudy (anonymous):

\[\text{or }12=2 \cdot(2 + 2/2)!.\]So complicated LOL. I can't think for 13.

OpenStudy (isaiah.feynman):

Me too

OpenStudy (anonymous):

@pgpilot326 said 13 = 22/2 + 2.

OpenStudy (anonymous):

\[ 6=\frac{(2+2)!}{2+2} \]

OpenStudy (anonymous):

\[2^{2^{2}}-2 = 14\]

OpenStudy (isaiah.feynman):

Nice one @pgpilot326

OpenStudy (anonymous):

also \[\frac{(2+2)!}{2}+2=14

OpenStudy (anonymous):

oops \[\frac{(2+2)!}{2}+2=14\]

OpenStudy (anonymous):

A medal for working for 13. :)

OpenStudy (isaiah.feynman):

I figured out the one for 15 but it used 5 2's.

OpenStudy (anonymous):

Answer to the spin-off question: \(2^{2^{22}}\). I don't know how it was obtained, but that's the answer I saw. :)

OpenStudy (anonymous):

\[\sqrt{22^2}-2=20\]

OpenStudy (isaiah.feynman):

Easy \[16=2^{3}\times2\]

OpenStudy (isaiah.feynman):

An alternative way for 16 is welcome.

OpenStudy (anonymous):

\[2^2\cdot 2^2 = 16\]

OpenStudy (anonymous):

\[(2+2)^2=16\]

OpenStudy (isaiah.feynman):

Even better using 3 two's.

OpenStudy (anonymous):

\[2^{2^{2}}=16\]

OpenStudy (isaiah.feynman):

\[15= (2+2)^{2}-2^{0}\]

OpenStudy (anonymous):

so you can use numbers besides 2?

OpenStudy (isaiah.feynman):

No. I broke the rule.

OpenStudy (anonymous):

\[\left(\begin{matrix}2+2+2 \\ 2\end{matrix}\right)=15\]

OpenStudy (anonymous):

17

OpenStudy (anonymous):

22-2-2=18

OpenStudy (anonymous):

\[(2\cdot 2+\frac{2}{2})!!=15\]

OpenStudy (anonymous):

\[(2+\frac{2}{2})^2=9\]

OpenStudy (anonymous):

Can you get 7?

OpenStudy (anonymous):

\[(2+2)!!-\frac{2}{2}\]

OpenStudy (anonymous):

also \[(2+2)!!+\frac{2}{2}=9\]

OpenStudy (anonymous):

What is !!?

OpenStudy (anonymous):

double factorial

OpenStudy (anonymous):

\[3!! = 1\cdot 3\] etc. \[4!!=2\cdot 4\] etc.

OpenStudy (anonymous):

0!!=1

OpenStudy (anonymous):

22/2 = 17 in hex

OpenStudy (anonymous):

doesn't have to be base 10, right?

OpenStudy (anonymous):

and then 22/2 + 2 = 19 in hex

OpenStudy (isaiah.feynman):

The only rule I know is that you must use a maximum of 4 2's

OpenStudy (anonymous):

winner, winner, chicken dinner!

OpenStudy (anonymous):

i'm gonna use this in my classes... what a great puzzle!

OpenStudy (anonymous):

22-2=20, 22-2/2=21, 22=22, 22+2/2=23, (2+2)!=24, (2+2)! + 2/2 = 25 22+2+2 = 26

OpenStudy (isaiah.feynman):

Are you a maths teacher?

OpenStudy (anonymous):

sometimes

OpenStudy (anonymous):

2(22-2)=28 in base 7 (2+2)! + 2 + 2 = 28

OpenStudy (whpalmer4):

another method for 16: \[(2+2)*(2+2)= 16\]

OpenStudy (anonymous):

\[2^{2^{2^{2}}}=2^{2^{4}}=2^{16}=65536\] Maybe throw in a factorial sign or two in there. \[2^{222!}\] or \[22!^{22!}!\] Both are incredibly larger than the first one I gave and can't be expressed by Google, apparently.

OpenStudy (the_fizicx99):

Fun ♦o♦